For which values of a and b is the following equation true?

lim (sinx/x³ + a + b/x²) = 0
x - 0
1
Expert's answer
2020-03-03T16:12:58-0500

The limit equation is given by:

"\lim_{x \to 0} ( \frac{\sin (x)}{x^3} + a + \frac{b}{x^2} ) = 0"

Separating the constant term "a" from the limit and taking limit for the other limit function, we have:

"\lim_{x \to 0 } ( \frac{\sin (x)}{x^3} + \frac{b}{x^2}) + a = 0"

Now:

The limit value of the limit function is:

"\lim_{x \to 0} ( \frac{ \sin(x) + bx}{x^3} )"

Inserting "x = 0" into the limit function, we can see that we have "\frac{0}{0}" form occurs. For this "\frac{0}{0}" form, we will use L'Hospital's Rule. Using this rule, we will differentiate both numerator and denominator of the limit function. After differentiating both numerator and denominator of the limit function, we have:

"\lim_{x \to 0} \left( \frac{ \cos (x) + b}{3x^2} \right)"

"= \frac{\cos(0) + b}{3(0^2)} \hspace {1 cm}" [ By substituting "x = 0" into the limit function ]

"= \frac{1 + b}{0} \hspace{1 cm}" [ "\because \cos (0) = 1" ]

"= \infty"

As, the right hand side of the limit equation is so, we must keep this limit value to . For this, "b + 1" must be equal to

So:

"b + 1 = 0 \\
\Rightarrow b = -1"

Now:

Inserting "b = -1" into the limit equation, we have:

"\lim_{x \to 0} ( \frac{\sin (x)}{x^3} - \frac{1}{x^2} ) + a = 0"

"\Rightarrow \lim_{x \to 0} ( \frac{\sin (x) - x}{x^3} ) + a = 0"

"\Rightarrow \lim_{x \to 0} ( \frac{ \cos (x) - 1 }{3x^2} ) + a = 0" [ "\because \frac{0}{0}" form occurs so, we use L'Hospital's Rule ]

"\Rightarrow \lim_{x \to 0} ( \frac{ - \sin (x)}{6x} ) + a = 0" [ Again, "\frac{0}{0}" form occurs so, we use L'Hospital's Rule ]

"\Rightarrow - \lim_{x \to 0} ( \frac{\cos (x)}{6} ) + a = 0" [ Again, "\frac{0}{0}" form occurs so, we use L'Hospital's Rule]

"\Rightarrow - \frac{\cos(0)}{6} + a= 0" [ Substituting "x = 0" into the limit function ]

"\Rightarrow - \frac{1}{6} + a = 0" [ "\because \cos (0) = 1" ]

"\Rightarrow a = \frac{1}{6}"

Hence:

The values of "a \,\, and \,\, b" are "\frac{1}{6} \, and \, -1" respectively.